OpenStudy (anonymous):

x^2-4y^2=9 y-2x=0, solve by method of substitution can someone show me the steps for this problem

OpenStudy (cwrw238):

y = 2x from the second equation now replace y in the first equation to get an equation in x only solve this equation for x and substitute these value(s) in first equation to find y

OpenStudy (cwrw238):

second line replace y by 2x in first equation ...

OpenStudy (anonymous):

do u think you can show me step by step

OpenStudy (anonymous):

sorry i hate math :(

OpenStudy (cwrw238):

gotta go sorry

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

x^2-4y^2=9-(1) y-2x=0 y=2x Substitute this in (1) x^2-4(2x)^2=9 x^2-4*4x^2=9 x^2-16x^2=9 -15x^2=9 x^2=-9/15 =-3/5 x=sqrt(-3/5) y=2*sqrt(-3/5) Hope it's clear. sorry gotta go nw.

OpenStudy (anonymous):

ok thanks for your help :)

OpenStudy (ajprincess):

yw.:)